# Learning noisy tissue dynamics across time scales

### v1

@@ Abstract
Tissue dynamics play a crucial role in biological processes ranging from wound healing to morphogenesis. However, these noisy multicellular dynamics are notoriously hard to predict. Here, we introduce a biomimetic machine learning framework capable of inferring noisy multicellular dynamics directly from experimental movies. This generative model combines graph neural networks, normalizing flows and WaveNet algorithms to represent tissues as neural stochastic differential equations where cells are edges of an evolving graph. This machine learning architecture reflects the architecture of the underlying biological tissues, substantially reducing the amount of data needed to train it compared to convolutional or fully-connected neural networks. Taking epithelial tissue experiments as a case study, we show that our model not only captures stochastic cell motion but also predicts the evolution of cell states in their division cycle. Finally, we demonstrate that our method can accurately generate the experimental dynamics of developmental systems, such as the fly wing, and cell signaling processes mediated by stochastic ERK waves, paving the way for its use as a digital twin in bioengineering and clinical contexts.
The physicist Eugene Wigner famously commented on the “unreasonable effectiveness of mathematics in the natural sciences” Wigner (1960). Continuum theories are an example of such unreasonable effectiveness Gaspard (2022). These deterministic, memoryless, coarse-grained theories describe, often with uncanny precision, systems that are discrete, stochastic, and non-Markovian. Yet, this unholy trinity of complexity comes back to haunt us when we seek top-down approaches that infer microscopic rules, including single-cell variability and noise, from the macroscopic behavior of biological systems. Inspired by Wigner’s creed, here we ask: What modeling framework, if any, can reliably predict the dynamics of noisy multicellular systems across time scales?
@@
Unlike the interactions between atoms or molecules, the precise biochemical mechanisms regulating cell dynamics remain largely unknown Phillips et al. (2013). This complexity arises from the intricate interplay of physical and chemical reactions among myriads of biological molecules inside a cell. Schematically, cell dynamics involve three key processes: motion, state transitions, and signaling (Fig. 1A). These processes challenge current algorithms, as they require to incorporate stochastic signals with unknown statistics and inherently discrete processes over multiple time scales. Moreover, cell interactions are both nonreciprocal and path dependent, since they can be mediated by complex cell signaling cascades. These difficulties notwithstanding, rapid advances in imaging routinely allow the automated collection of tissue dynamics datasets with single-cell resolution. This raises the prospect of developing machine learning methods for the study of multi-cellular dynamics that infer discrete probabilistic models from videos of tissues Rafelski and Theriot (2024); Schoenholz et al. (2016a); Berthier and Reichman (2023); Brückner and Broedersz (2024); Howe and Mani (2025); Cubuk et al. (2015); Sharp et al. (2018); Boattini et al. (2020); Schmitt et al. (2024); Brandstätter et al. (2025); Joshi et al. (2022); Brückner et al. (2021); Gkeka et al. (2020); Schoenholz et al. (2016b); Frishman and Keren (2021); Bapst et al. (2020); Yang et al. (2024); Yamamoto et al. (2022). In order to do so, we have to face the three challenges intrinsic to biological tissues summarized in Fig. 1B-C: discreteness, stochasticity, and non-Markovianity.
@@
First, a natural candidate to describe tissues are graph neural networks (GNN), a variant of neural networks  Scarselli et al. (2009); Zhou et al. (2020) capable of handling discrete data on irregular graphs of varying connectivity through message-passing algorithms that transmit information along their edges. In a nutshell, the architecture of these neural networks reflects the underlying biology of tissues: cells are represented by nodes in a dynamic graph, and cell signaling is mimicked by the message-passing between the nodes (Fig. 1A). Crucially, the GNN architecture encodes the fact that the units (cells) are indistinguishable from each other through permutation equivariance, and additionally incorporates an implicit bias towards local correlations. The main practical advantage of this biomimetic architecture is to substantially reduce the amount of data required to train compared to convolutional or fully connected deep neural networks that would have to rediscover this locality from data. This requirement can in principle be solved by having more samples, but this is impractical due to the intrinsic variability of biological systems. As the cell state is encoded in the nodes, the isotropic graph convolution underlying standard GNNs is not suitable to capture directed cascades of signaling events that involve correlations between subsequent signaling steps. In order to handle these, we augment the cell graph with a dual signaling graph where interactions are encoded in nodes (see Fig. 1A and Methods). It turns out that the combination of the cell graph and the signaling graph is sufficient to handle signaling cascades.
@@
Second, the stochastic dynamics observed in tissues requires modeling probability distributions and transition probabilities. In order to do so, we combine GNNs with normalizing flows, a class of generative models used to generate and sample complex probability distributions Tabak and Vanden-Eijnden (2010); Rezende and Mohamed (2015); Kobyzev et al. (2021). One of the key deliverables of this approach is the inference of individual cell-level probability distributions from a single tissue-level experiment. The reason why this is useful is the following: while it is known that noise can have important implications in biological systems Amir and Balaban (2018); Golding and Amir (2024); Höfling and Franosch (2013); Paulsson (2005); Tsimring (2014), it is not easy to perform reliable measurements of individual cell level stochasticity within tissues due to the high variability of biological samples.
@@
Third, the dynamics of biological tissues is often non-Markovian, an inescapable consequence of the fact that we cannot explicitly account for all the processes at play within each cell. Hence, one needs to keep track of the tissue over multiple time scales in order to predict its future. This is particularly challenging when these time scales are very different, because it requires handling an enormous amount of data all at the same time. In order to tackle this challenge, we take inspiration from WaveNet, a generative model recently developed to generate natural-sounding speech that mimics human voices van den Oord et al. (2016). In the case of voice, for instance, the duration of a spoken word is of the order of seconds, while the pitch of sounds can go from 100 Hz to 1000 Hz. Crucially, however, it is not necessary to resolve the whole audio signal at the shortest time scales. It suffices to describe the slowly-varying modulation of a carrier signal, like in AM radio. In a nutshell, WaveNet combines a version of this multiple time-scale approach, known as a causal dilated convolution, with autoregressive generative models. This strategy applied to GNNs yields graph WaveNets Wu et al. (2019).
Going back to biological tissues, using graph WaveNets in conjunction with normalizing flows, we can accurately generate the microscopic (e.g. single cell) dynamics, crucially including biological variability and noise, starting from experimental measurements of a multicellular system. This can, for instance, be used as a building block for creating digital twins of living tissues, a much sought-after goal for clinical and bioengineering applications Bunne et al. (2024); Katsoulakis et al. (2024).
@@
It describes the state ([Uncaptioned image]) of each cell ([Uncaptioned image]) by a set of continuous random variables that can include their position, velocity, and the concentration of certain proteins and genes expressed by the cells globally. The peculiar feature of this neural equation is that both terms on the right hand side are neural networks. The neural stochastic differential equation (1) is composed of a deterministic drift [Uncaptioned image], which may arise from a developmental cell fate, and a fluctuating noise [Uncaptioned image] which needs not be Gaussian because it could be generated by the internal active dynamics of the cell. For instance, in the dynamics of fate decision during cell differentiation, the drift term would represent the gene regulatory network capturing the complex dynamics among numerous transcription factors Karlebach and Shamir (2008). The noise term would capture the intrinsic stochasticity of chemical reactions in the process of gene expression and DNA binding events of transcription factors.
@@
Figure 2: Learning cell dynamics. A. We aim at predicting the velocity $\bm{v}_{i}$ of all individual cells from their positions $\bm{r}_{i}$ and cell shapes and areas ($c_{i}$) in tissues. This requires handling discrete and stochastic data, but no non-Markovian effect. B-F. We apply the technique to a monolayer of epithelial MDCK cells, which undergo stochastic migration and division in a confined environment (panel B). Cells become more rounded as they divide and gradually transit from a fluid state into a glassy state. Their motility, measured by the spatially averaged cell speed, decreases over time. Our ML model not only captures the same decreasing trend (panel F), but also accurately predicts the spatial variation (shaded region). Panel C shows the velocity field at single-cell resolution. Even though an experimental movie only provides a particular realization of the stochastic dynamics, the probabilistic design enables us to infer the underlying statistics through maximum likelihood optimization. This allows us to further extract the ensemble properties of individual cells (panel E). G-J. We applied our ML algorithm to study the growth of a fly wing Etournay et al. (2015, 2016). It automatically recognizes the deterministic nature of such a developmental system. The predicted probability distribution of cell migration velocity is nearly a $\delta$-function (panel J). Note that all the ML predictions shown in this paper are produced on the test dataset, which has never been seen by the ML model during training. K The ML pipeline allows us to predict the evolution of both average features (red and blue lines) and noise (manifested through observed variability; red and blue regions). L. We aim at predicting the state of individual cells within the cell division cycle from the individual positions in the tissue. In the division process, the cells go through three states, $G_{1}$, $S$, and $G_{2}$ and then undergo mitosis ($M$). It is a stochastic process discrete in time meanwhile subject to regulations from neighboring cells through mechanical stresses, but here we do not consider non-Markovian effects. M. In a monolayer of MDCK cells, we can identify the division states of individual cells at any given time. N-O. Comparison between experimental ground truth (panel N) and ML predictions (panel O). After training, our model can predict the cell division state with accuracy over 80%.
@@
Stochastic effects, our second challenge (second row in Fig. 1B-C), are captured by the noise term [Uncaptioned image] in Eq. (1). In particular, we must allow the noise in the states of individual cells to be correlated with each other through the interaction network. In order to do so, we combine the graph neural networks (which handle deterministic rules) with normalizing flows, a class of generative models used to generate and sample complex probability distributions Tabak and Vanden-Eijnden (2010); Rezende and Mohamed (2015); Kobyzev et al. (2021). The resulting graph normalizing flow Liu et al. (2019) allows us to model the joint probability distribution of all the cells in the tissue. We enforce that different cells should behave identically, with the same drift term and the same probability distribution of the noise, when they have the same biological state [Uncaptioned image] and the same local environment (i.e. other cells in the graph [Uncaptioned image]), like real cells tend to do. It is this biomimetic architecture that allows us to infer cell-level probability distributions (that can be then sampled to compute averages and correlations) despite a reduced amount of data that would otherwise not be sufficient to reliably estimate these quantities.
@@
To capture memory effects, arising from our third challenge non-Markovianity (third row in Fig. 1B-C), we allow [Uncaptioned image] to be a function of the system history (represented by the time sequence [Uncaptioned image]) which describes how the states and interactions networks of each cell (represented by the graph [Uncaptioned image]) vary over time. Similarly, the noise is sampled from a joint probability distribution (represented by the symbol [Uncaptioned image]), which is also history dependent and allows correlations between cells that are generated through their interaction network. We encode long-term memory effects through causal dilated convolution layers, which perform multi-scale context aggregation, following the approach used in WaveNet van den Oord et al. (2016); Yu and Koltun (2015); Dutilleux (1990); Holschneider et al. (1989).
@@
We now show that our algorithmic framework allows us to efficiently learn both terms [Uncaptioned image] and [Uncaptioned image] of Eq. (1) at the same time from experimental movies. We present our examples in order of increasing computational difficulty starting from situations where the random term [Uncaptioned image] dominates (epithelial tissues, Fig. 2A-F) or where the deterministic drift [Uncaptioned image] dominates (development, Fig. 2G-J) culminating in the most challenging case of mixed stochastic and deterministic dynamics (cell signaling, Fig. 3).
@@
Predicting cell dynamics. To test our algorithmic framework, we proceed to predict the motion of cells solely from their geometry and position within the tissues. In this case, the deterministic drift term [Uncaptioned image] in Eq. (1) is negligible and the dynamics is governed by the fluctuating noise [Uncaptioned image] that is approximately Markovian, making the addition of the WaveNet part unnecessary in this simpler example. Prior models and experiments have demonstrated a relationship between cell geometry and tissue flow but only for average cell motions Bi et al. (2015, 2016); Park et al. (2015); Mongera et al. (2018); Petridou et al. (2021); Devany et al. (2021).
@@
Instead we task our GNNs with predicting the motion of each cell in the tissue at every instant (Fig. 2A). One of the main challenges of predicting the cell motion is the inherent stochasticity of motion that comes from active forces generated by each cell. Such forces can be ultimately traced to processes in the cytoskeleton and are not easily experimentally accessible. To account for such stochastic processes, we predict a probability distribution for the displacement of each cell based on the current configuration. Training is done for each system of interest which allows for the neural network to determine the level of stochasticity exhibited by cells in the tissue.
@@
We apply our algorithms to epithelial tissue mono-layers of Madin-Darby canine kidney (MDCK) cells, a well studied model system for collective cell migration Aoki et al. (2017); Rørth (2009); Hakim and Silberzan (2017) for which large amounts of data are available for training (Fig. 2B-F). We performed time lapse imaging experiments taking images of cell membranes and nuclei every 10 minutes over roughly 24 hours (Fig. 2B-C), see Methods for details. The cell membrane and nuclei were then segmented and tracked to produce a list of cell positions, sizes, shapes and displacements to input into the GNN. As it is typically not possible to correctly segment and track every cell, we have developed an automated method for filling in missing cells with predicted "dummy cells" that have the correct average properties. Figure 2C shows an example of the resulting data: cell centers are shown as points colored by speed and their velocities are represented by arrows. We train the network with all these inputs and task it to make a prediction of the displacement from the other geometric variables.
@@
As in previous works Devany et al. (2021), we demonstrate that the information about cell geometry correlates with the average cell speed in the monolayer (Fig. 2C-D). Crucially, the GNN also predicts the distribution of cell speeds at each time point as evidenced by the correspondence between the standard deviations (compare blue and red curves in Fig. 2F). Observing the displacements of each individual cell, we notice that the GNN is able to predict the small scale coherent flows seen in cell monolayers (Fig. 2C-D) Bi et al. (2016). Furthermore, we see the stochastic behavior of cells by the various cell displacements predicted by the GNN. Here, the predicted probability distribution (Fig. 2E) has multiple dominant peaks suggesting that for the current configuration, either of these displacements is favorable depending on forces applied by individual cells in the tissue. This level of stochasticity in the cell dynamics is in line with previous analysis showing that the displacements of cells do not correlate over long spatial or temporal scales Devany et al. (2021).
@@
We now move on to a different example where we compare and contrast the results presented above with predictions made on a developing fly wing (Fig. 2G-J). To do so, we applied our GNN analysis to published fly wing data Etournay et al. (2015). In this case, the deterministic drift term [Uncaptioned image] in Eq. 1 dominates and it is the noise [Uncaptioned image] that is small. While predictive models at the cell scale for such systems are still lacking, developmental systems show more stereotypical behavior across embryos amenable to continuum modeling approaches Streichan et al. (2018); Lefebvre et al. (2023); Hannezo and Heisenberg (2019); Garcia et al. (2015); Saadaoui et al. (2020); Streichan et al. (2018); Tavano et al. (2025); Brückner and Hannezo (2024). Our algorithm successfully predicts motion across the wing despite only being trained on a subset of a single experiment Etournay et al. (2015, 2016) (compare Fig. 2H and I). We observe that the predicted distribution of displacements are noticeably more deterministic than the monolayers in vitro showing a single strong peak in the probability distribution (Fig. 2J, compare with the in vitro monolayer in panel E). This matches our intuition that the developmental system should behave in a reproducible manner while the MDCK monolayer may be much more chaotic. We emphasize that the inferred single-cell probability distributions in Fig. 2E and J correspond to a single experiment (plus an already-trained network). This ability to infer ensemble quantities from a single realization is a key strength of our approach.
@@
Predicting cell state transitions. Next, we attempt to predict the internal biological state of the cell from cell geometry only. We focus on the cell cycle, the sequence of proliferation and division (Fig. 2L). The cell cycle is comprised into four phases (Fig. 2L): the $G_{1}$ phase, a growth phase prior to DNA replication; the S phase, when DNA is replicated, and the $G_{2}$ phase, a second growth phase when the cell prepares to enter mitosis (M). The $G_{1}$ phase has the largest variation in duration, which is determined by whether conditions are met to proceed to S phase. By contrast, in MDCK cells, the S and $G_{2}$ phases are typically 8-10 and 2-4 hours, respectively. In epithelial monolayers geometric variables are known to correlate with the cell cycle. For example, the number of neighbors a cell has depends on the relative time since dividing compared to other cells Gibson et al. (2006); Gibson and Gibson (2009), and the area of a cell can be used to infer if it will become cell cycle arrested due to contact inhibition Devany et al. (2023). However, a comprehensive model which can make accurate predictions at the single cell level does not exist. We asked if the GNN could integrate these geometric information to make an accurate prediction of the cell state. This is challenging because of the stochastic nature of the cell state transitions between phases of the cell cycle, the discreteness of signaling states and the relationship between the properties of neighboring cells.

### v2

@@ Abstract
Tissue dynamics play a crucial role in biological processes ranging from inflammation to morphogenesis. However, these noisy multicellular dynamics are notoriously hard to predict. Here, we introduce a biomimetic machine learning framework capable of inferring noisy multicellular dynamics directly from experimental movies. This generative model combines graph neural networks, normalizing flows and WaveNet algorithms to represent tissues as neural stochastic differential equations where cells are edges of an evolving graph. Cell interactions are encoded in a dual signaling graph capable of handling signaling cascades. The dual graph architecture of our neural networks reflects the architecture of the underlying biological tissues, substantially reducing the amount of data needed for training, compared to convolutional or fully-connected neural networks. Taking epithelial tissue experiments as a case study, we show that our model not only captures stochastic cell motion but also predicts the evolution of cell states in their division cycle. Finally, we demonstrate that our method can accurately generate the experimental dynamics of developmental systems, such as the fly wing, and cell signaling processes mediated by stochastic ERK waves, paving the way for its use as a digital twin in bioengineering and clinical contexts.
The physicist Eugene Wigner famously commented on the “unreasonable effectiveness of mathematics in the natural sciences” Wigner (1960). Continuum theories are an example of such unreasonable effectiveness Gaspard (2022). These deterministic, memoryless, coarse-grained theories describe, often with uncanny precision, systems that are discrete, stochastic, and non-Markovian. Yet, this triad of complexity comes back to haunt us when we seek top-down approaches that infer microscopic rules, including single-cell variability and noise, from the macroscopic behavior of biological systems. Inspired by Wigner’s creed, here we ask: What modeling framework, if any, can reliably predict the dynamics of noisy multicellular systems across time scales?
@@
Unlike the interactions between atoms or molecules, the precise biochemical mechanisms regulating cell dynamics remain largely unknown despite their crucial role in biological processes ranging from morphogenesis to inflammation Phillips et al. (2013). This complexity ultimately arises from the intricate interplay of physical and chemical reactions among myriads of biological molecules inside a cell. Schematically, cell dynamics involve three key processes: motion, state transitions, and signaling (Fig. 1A). These processes challenge current algorithms, as they require to incorporate stochastic signals with unknown statistics and inherently discrete processes over multiple time scales. Moreover, cell interactions are both nonreciprocal and path dependent, since they can be mediated by complex cell signaling cascades. These difficulties notwithstanding, rapid advances in imaging routinely allow the automated collection of tissue dynamics datasets with single-cell resolution. This raises the prospect of developing machine learning methods for the study of multi-cellular dynamics that infer discrete probabilistic models of tissues Rafelski and Theriot (2024); Schoenholz et al. (2016a); Berthier and Reichman (2023); Brückner and Broedersz (2024); Howe and Mani (2025); Cubuk et al. (2015); Sharp et al. (2018); Boattini et al. (2020); Schmitt et al. (2024); Brandstätter et al. (2025); Joshi et al. (2022); Brückner et al. (2021); Gkeka et al. (2020); Schoenholz et al. (2016b); Frishman and Keren (2021); Bapst et al. (2020); Yang et al. (2024); Yamamoto et al. (2022). In order to do so, we face three challenges intrinsic to biological tissues illustrated in Fig. 1B-C: discreteness, stochasticity, and non-Markovianity.
@@
First, a natural candidate to describe tissues are graph neural networks (GNN), a variant of neural networks  Kipf and Welling (2017); Scarselli et al. (2009); Zhou et al. (2020) capable of handling discrete data on irregular graphs of varying connectivity through message-passing algorithms that transmit information along their edges. In a nutshell, the architecture of these neural networks reflects the underlying biology of tissues: cells are represented by nodes in a dynamic graph, and cell signaling is mimicked by the message-passing between the nodes (Fig. 1A). Crucially, the GNN architecture encodes the fact that the units (cells) are indistinguishable from each other through permutation equivariance, and additionally incorporates an implicit bias towards local correlations. The main practical advantage of this architecture is to substantially reduce the amount of data required to train compared to convolutional or fully connected deep neural networks that would have to rediscover locality from data. This requirement can in principle be solved by having more samples, but this is impractical for tissues due to the intrinsic variability of these biological systems. As the cell state is encoded in the nodes, the isotropic graph convolution underlying standard GNNs is not suitable to capture directed cascades of signaling events that involve correlations between subsequent signaling steps. The key methodological innovation that allows us to handle this crucial biological complexity is to augment the cell graph with a dual signaling graph where interactions are encoded in nodes (see Fig. 1A and Methods). It is this biomimetic architecture of our dual-graph NNs that ultimately captures signaling cascades.
@@
Second, the stochastic dynamics observed in tissues requires modeling probability distributions and transition probabilities. In order to do so, we combine GNNs with normalizing flows, a class of generative models that can be used to efficiently generate, manipulate, and sample complex probability distributions Tabak and Vanden-Eijnden (2010); Rezende and Mohamed (2015); Kobyzev et al. (2021); Liu et al. (2019). One of the key deliverables of this approach is the inference of individual cell-level probability distributions from a single tissue-level experiment. The reason why this is useful is the following: while it is known that noise can have important implications in biological systems Amir and Balaban (2018); Golding and Amir (2024); Höfling and Franosch (2013); Paulsson (2005); Tsimring (2014), it is not easy to perform reliable measurements of individual cell level stochasticity within tissues.
@@
Third, the dynamics of biological tissues is often non-Markovian, an inescapable consequence of the fact that we cannot explicitly account for all the processes at play within each cell. Hence, one needs to keep track of the tissue over multiple time scales in order to predict its future. This is particularly challenging when these time scales are very different, because it requires handling large amounts of data all at once. In order to tackle this challenge, we take inspiration from WaveNet, a generative model developed to generate natural-sounding speech that mimics human voices van den Oord et al. (2016); Wu et al. (2019). In the case of voice, for instance, the duration of a spoken word is of the order of seconds, while the pitch of sounds can go from 100 Hz to 1000 Hz. Crucially, however, it is not necessary to resolve the whole audio signal at the shortest time scales. It suffices to describe the slowly-varying modulation of a carrier signal, like in AM radio. In a nutshell, WaveNet combines a version of this multiple time-scale approach, known as a causal dilated convolution, with autoregressive generative models. This strategy applied to GNNs is henceforth referred to as graph WaveNets.
We now go back to biological tissues and provide a step by step guide on how to combine dual-graph WaveNets and normalizing flows into a coherent modeling framework, using the three challenges as organizing principles of our exposition.
@@
It describes the state ([Uncaptioned image]) of each cell ([Uncaptioned image]) by a set of continuous random variables that can include their position, velocity, and the concentration of certain proteins and genes expressed by the cells globally. What makes Eq. (1) a neural equation is that both terms on the right hand side are neural networks. The neural stochastic differential equation (1) is composed of a deterministic drift [Uncaptioned image], which may arise from a developmental cell fate, and a fluctuating noise [Uncaptioned image] which needs not be Gaussian because it could be generated by the internal active dynamics of the cell. For instance, in the dynamics of fate decision during cell differentiation, the drift term would represent the gene regulatory network capturing the complex dynamics among numerous transcription factors Karlebach and Shamir (2008). The noise term would capture the intrinsic stochasticity of chemical reactions in the process of gene expression and DNA binding events of transcription factors.
@@
Figure 2: Learning cell dynamics. A. We aim at predicting the velocity $\bm{v}_{i}$ of all individual cells from their positions $\bm{r}_{i}$ and cell shapes and areas ($c_{i}$) in tissues. This requires handling discrete and stochastic data, but no non-Markovian effect. B. The ML pipeline allows us to predict the evolution of both average features (red and blue lines) and noise (manifested through observed variability; red and blue regions). C-F. We apply the technique to a monolayer of epithelial MDCK cells, which undergo stochastic migration and division in a confined environment (panel C). Cells become more rounded as they divide and gradually transit from a fluid state into a glassy state. Their motility, measured by the spatially averaged cell speed, decreases over time. Our ML model not only captures the same decreasing trend (panel B), but also accurately predicts the spatial variation (shaded region). Panel D shows the velocity field at single-cell resolution. Even though an experimental movie only provides a particular realization of the stochastic dynamics, the probabilistic design enables us to infer the underlying statistics through maximum likelihood optimization (panel E). This allows us to further extract the ensemble properties of individual cells (panel F). G-J. We applied our ML algorithm to study the growth of a fly wing Etournay et al. (2015, 2016). It automatically recognizes the deterministic nature of such a developmental system. The predicted probability distribution of cell migration velocity is nearly a $\delta$-function (panel J). Note that all the ML predictions shown in this paper are produced on the test dataset, which has never been seen by the ML model during training. L. We aim at predicting the state of individual cells within the cell division cycle from the individual positions in the tissue. In the division process, the cells go through three states, $G_{1}$, $S$, and $G_{2}$ and then undergo mitosis ($M$). It is a stochastic process discrete in time meanwhile subject to regulations from neighboring cells through mechanical stresses, but here we do not consider non-Markovian effects. M. In a monolayer of MDCK cells, we can identify the division states of individual cells at any given time. N-O. Comparison between experimental ground truth (panel N) and ML predictions (panel O). After training, our model can predict the cell division state with accuracy over 80%.
@@
Stochastic effects, our second challenge (second row in Fig. 1B-C), are captured by the noise term [Uncaptioned image] in Eq. (1). In particular, we must allow the noise in the states of individual cells to be correlated with each other through the interaction network. In order to model the joint probability distribution of all the cells in the tissue, we resort to our dual-graph normalizing flow architecture. We enforce that different cells should behave identically, with the same drift term and the same probability distribution of the noise, when they have the same biological state [Uncaptioned image] and the same local environment (i.e. other cells in the graph [Uncaptioned image]), like real cells tend to do. The key deliverable of our biomimetic algorithms is that they are capable of inferring single-cell-level probability distributions (that can be then sampled to compute averages and correlations), despite using only a reduced amount of data that would otherwise not suffice to reliably estimate these quantities.
@@
To capture memory effects, arising from our third challenge non-Markovianity (third row in Fig. 1B-C), we allow [Uncaptioned image] to be a function of the system history (represented by the time sequence [Uncaptioned image]) which describes how the states and interactions networks of each cell (represented by the graph [Uncaptioned image]) vary over time. Similarly, the noise is sampled from a joint probability distribution (represented by the symbol [Uncaptioned image]), which is also history dependent and allows correlations between cells generated through their interaction network. We encode long-term memory effects through causal dilated convolution layers, which perform multi-scale context aggregation, following the approach used in WaveNet van den Oord et al. (2016); Yu and Koltun (2015); Dutilleux (1990); Holschneider et al. (1989).
@@
We now show that our algorithmic framework allows us to efficiently learn both terms [Uncaptioned image] and [Uncaptioned image] of Eq. (1) at the same time from experimental movies. We present our examples in order of increasing computational difficulty starting from situations where the random term [Uncaptioned image] dominates (epithelial tissues, Fig. 2A-F) or where the deterministic drift [Uncaptioned image] dominates (development, Fig. 2G-J) culminating in the most challenging case of mixed stochastic and deterministic dynamics (cell signaling mediated by ERK waves, Fig. 3).
@@
Noisy cell dynamics in epithelia. As a first application of our algorithmic framework, we proceed to tackle a basic biophysical question that is still open: is the noisy motion of cells determined (i.e. predictable) solely by their geometry and position within the tissue? In this case, the deterministic drift term [Uncaptioned image] in Eq. (1) is negligible and the dynamics is dominated by the fluctuating noise [Uncaptioned image] that is approximately Markovian, making the addition of the WaveNet part unnecessary in this simpler example. Prior models and experiments have suggested that a correlation exist between cell geometry and tissue flow but only for average cell motions Bi et al. (2015, 2016); Park et al. (2015); Mongera et al. (2018); Petridou et al. (2021); Devany et al. (2021).
@@
Instead we task our GNNs with predicting the stochastic dynamics of each cell in the tissue at every instant (Fig. 2A). What makes this noisy cell motion hard to predict is that it is generated by intracellular active forces ultimately traceable to stochastic processes within the cytoskeleton not easily accessible experimentally. This biohysical hurdles notwithstanding, we can predict a probability distribution for the displacement of each cell based on the current configuration. Training is done for each system of interest which allows for the neural network to determine the level of stochasticity exhibited by cells in the tissue.
@@
We apply our algorithms to epithelial tissue mono-layers of Madin-Darby canine kidney (MDCK) cells, a well studied model system for collective cell migration Aoki et al. (2017); Rørth (2009); Hakim and Silberzan (2017) for which large amounts of data are available for training (Fig. 2B-F). We performed time lapse imaging experiments taking images of cell membranes and nuclei every 10 minutes over roughly 24 hours (Fig. 2C-D), see Methods for details. The cell membrane and nuclei were then segmented and tracked to produce a list of cell positions, sizes, shapes and displacements to input into the GNN. As it is typically not possible to correctly segment and track every cell, we have developed an automated method for filling in missing cells with predicted "dummy cells" that have the correct average properties. Figure 2D shows an example of the resulting data: cell centers are shown as points colored by speed and their velocities are represented by arrows. We train the network with all these inputs and task it to make a prediction of the displacement from the other geometric variables.
@@
We now attempt to discover the full probability distribution of cell speeds using our generative model even if a direct measurement of this quantity is not possible in our small dataset without the implicit data augmentation enabled by our equivariant ML architecture (Fig. 2A-C). Inspection of Fig. 2B reveals that at each time point, the prediction (red) made by our neural networks fully matches the experimental data (blue), not only for the mean (continuous line) but also the standard deviation (shaded area) of cell speed. We emphasize that this prediction is made only from cells geometry and positions. Observing the displacements of each individual cell, we notice that the GNN is able to predict the small scale coherent flows seen in cell monolayers (Fig. 2D-E) Bi et al. (2016). Furthermore, we see the stochastic behavior of cells by the various cell displacements predicted by the GNN. Here, the predicted probability distribution (Fig. 2F) has multiple dominant peaks suggesting that for the current configuration, either of these displacements is favorable depending on forces applied by individual cells within tissues.
@@
Deterministic cell dynamics in morphogenesis. We now move on to a different example where we compare and contrast the results presented above with predictions made on a developing fly wing (Fig. 2G-J). To do so, we applied our GNN analysis to published fly wing data Etournay et al. (2015). In this case, the deterministic drift term [Uncaptioned image] in Eq. (1) dominates and it is the noise [Uncaptioned image] that is small. While predictive models at the cell scale for such systems are still lacking, developmental systems show more stereotypical behavior across embryos amenable to continuum modeling approaches Streichan et al. (2018); Lefebvre et al. (2023); Hannezo and Heisenberg (2019); Garcia et al. (2015); Saadaoui et al. (2020); Streichan et al. (2018); Tavano et al. (2025); Brückner and Hannezo (2024). Our algorithm successfully predicts motion across the wing despite only being trained on a subset of a single experiment Etournay et al. (2015, 2016) (compare Fig. 2H and I). We observe that the predicted distribution of displacements are noticeably more deterministic than the monolayers in vitro showing a single strong peak in the probability distribution (Fig. 2J, compare with the in vitro monolayer in panel E). This matches our intuition that the developmental system should behave in a reproducible manner while the MDCK monolayer may be much more chaotic. We emphasize that the inferred single-cell probability distributions in Fig. 2F and J correspond to a single experiment (plus an already-trained network). This ability to infer ensemble quantities from a single realization ilustrates a key practical strength of our approach.
@@
Cell state transitions in the cell cycle. Next, we attempt to predict the internal biological state of the cell from cell geometry only. We focus on the cell cycle, a sequence of proliferation and division composed of four phases (Fig. 2L): the $G_{1}$ phase, a growth phase prior to DNA replication; the $S$ phase, when DNA is replicated, and the $G_{2}$ phase, a second growth phase when the cell prepares to enter mitosis ($M$). The $G_{1}$ phase has the largest variation in duration, which is determined by whether conditions are met to proceed to the $S$ phase. By contrast, in MDCK cells, the $S$ and $G_{2}$ phases are typically 8–10 and 2–4 hours, respectively. In epithelial monolayers geometric variables are known to correlate with the cell cycle. For example, the number of neighbors a cell has depends on the relative time since dividing compared to other cells Gibson et al. (2006); Gibson and Gibson (2009), and the area of a cell can be used to infer if it will become cell cycle arrested due to contact inhibition Devany et al. (2023). However, a comprehensive model which can make accurate predictions at the single cell level does not exist. We asked the GNN to integrate these geometric information to make a prediction of the cell state. This is challenging because of the stochastic nature of the cell state transitions between phases of the cell cycle, the discreteness of signaling states and the relationship between the properties of neighboring cells.